# Advancing Galaxy Analysis: AI-Powered Detection and Segmentation of Edge-On Galaxies

### v1

@@ 3.1 YOLOv5 Training and Evaluation
For the detection task, we used a manually annotated set of images from SDSS. The annotation was performed by the students using the Zooniverse platform. The results of the performance of the YOLOv5 model on the test set are summarised in Table 1. The table shows the number of images and its instances of edge-on galaxies, along with the key metrics: Precision and Recall. These results indicate that the model performs well in detecting edge-on galaxies, with a balanced trade-off between precision and recall.
@@
Table 1: Table with metrics calculated by YOLOv5 on testing dataset. One image may contain more than 1 instance of edge-on galaxy. Precision and Recall are calculated according to TP, FP and FN values (see Table 2).
@@
To provide a comprehensive understanding of the performance of the YOLOv5 model, we also provide particular detection cases, where we can see the sources of errors more precisely. Table 2 summarises the counts of TP, FP, and FN, which are essential for estimating the detection capabilities of the model.
Table 2: Summary of TPs, FPs, and FNs identified by the YOLOv5 detection model.
@@
Due to manual annotations by students, there is some inevitable annotation bias, e.g., not every time students annotated all instances or sometimes a wrong object was annotated. One of the advantages of deep learning techniques is their ability to overcome annotation bias issues and provide correct predictions even for cases with annotation mistakes. This is possible if annotations are generally correct for most cases and the architecture is able to learn a good generalised prediction function. Hence, the real efficiency of algorithm is even better and might be additionally estimated by the analysis of wrong cases.
According to our knowledge from the Galaxy Zoo project, there should be at least one edge-on galaxy in the image, but that does not rule out the possibility of detecting other galaxies. Initially, our annotators only flagged galaxies that they were confident of, but our algorithm successfully detected several others. This is illustrated in Fig. 2, where we present an example image that initially had one galaxy annotation from Zooniverse (see Fig. 1). After applying our YOLOv5 model, an additional galaxy was detected. This demonstrates the capability of the model to identify galaxies that may have been missed in the initial annotations due to manual annotation bias.
@@
To verify these findings, we manually examined detections labeled as FP and FN. Figures 3 and 4 show cases of FPs and FNs, providing visual context for these detections.
@@
Through visual inspection, we found that a number of misclassifications are due to inaccuracies in the annotation. In more detail, at least $\sim 20\%$ of the FP cases are considered edge-on galaxies, which were not annotated by mistake, and therefore these cases might be relocated from FP to TP after such analysis. For another $\sim 20\%$ of galaxies we were not able to decide, based on visual inspection, what their inclination is, but some of them might be edge-on as well. Hence, we can estimate that Precision is probably at least $\sim 85\%$ in comparison to $\sim 80\%$ measured for the original annotations. Although these numbers are only approximate, as they were obtained by visual inspection, which is prone to bias, they suggest that performance indicators may be underestimated and could potentially be higher with more accurate initial data. This observation highlights one of the key strengths of deep learning models mentioned before: the ability to learn the correct prediction function, even in the presence of partially imperfect annotations.
3.2 Segmentation of galaxies based on U-Net
Our main motivation behind this work is the analysis of warps in edge-on galaxies, which we plan to present in a future work. Therefore, to analyse our detected galaxies, we need to segment the image and extract the galaxies, while removing nearby objects that might contaminate them. There are several methods of image segmentation, such as thresholding, when the object is selected based on having a significantly higher pixel intensity value than the background (e.g., Lotz et al., 2004; Zee et al., 2022), using isophotes to define the extent of the galaxy (e.g., Hao et al., 2006; Reshetnikov et al., 2016), deep learning (e.g., Boucaud et al., 2020; Farias et al., 2020) and other methods (for a comprehensive list of methods and their comparisons see Xu & Zhu, 2024). Each method has its advantages and pitfalls, but in all cases, determining the threshold where the galaxy ends and the background starts exactly can be challenging.
We decided to extract the galaxies using our own segmentation algorithm. We used the bounding box coordinates obtained from the YOLOv5 detection to accurately crop galaxies from the original images. This procedure ensured that the segmented images contained only relevant galaxy data, thus optimising the segmentation process.
The process of creating high-quality segmentation masks from cropped galaxy images was time-consuming, mainly due to the manual adjustment of parameters. Each iteration involved careful parameter settings for thresholding and morphological operations such as erosion and dilation. After setting these parameters, we manually checked the resulting masks to ensure that they met our standards for clarity and definition of galaxy structures.
To automate the generation of segmentation masks, we used SCSS-Net, a convolutional neural network inspired by the U-Net architecture, originally designed to segment solar corona structures (Mackovjak et al., 2021). This versatile model was adapted for our specific requirement and retrained for segmentation of edge-on galaxies.
U-Net is a highly efficient convolutional neural network originally developed for medical image segmentation. Its architecture, designed to work effectively even with a limited amount of data, is characterised by its distinctive U-shape, comprising two main components: an encoder and a decoder (Ronneberger et al., 2015).
Encoder: The encoder is responsible for progressively capturing the context in the image, which involves reducing the spatial dimensions while increasing the depth. This process helps in extracting and compressing feature information, which is crucial for understanding the complex structures within the image.
Decoder: In contrast, the decoder part of the network works on upscaling the feature map to reconstruct the image output. It progressively recovers the spatial dimensions and detail of the image, ensuring that the important features extracted by the encoder are used to generate a precise and detailed segmentation. In our case the input is the original galaxy image with our manually created mask and the output is the mask created by the algorithm.
This architecture not only allows detailed image segmentation but also ensures that the network learns from a relatively small amount of data, making it particularly powerful in fields where high-quality data is scarce.
For the segmentation model learning, we divided our dataset into training, validation, and test sets, and tuned basic hyperparameters such as image size, batch size, and number of epochs. Due to the small size of our images ($64\times 64$ pixels), the initial training phase was efficient. We aimed for $1,000$ epochs with a batch size of $20$, but concluded training at epoch $963$ as performance gains plateaued.
  • Intersection over Union (IoU): Also known as the Jaccard Index, IoU is a statistical tool used to quantify the percent overlap between the target mask and the prediction output by the model. It is calculated as the area of overlap between the predicted segmentation and the ground truth divided by the area of union between the predicted and ground truth. This metric is particularly useful for determining how well the predicted boundaries align with the true boundaries.
    IoU=Area of OverlapArea of $\text{IoU}=\frac{\text{Area of Overlap}}{\text{Area of Union}}$ (5)
    Dice=2×|PredictionGround Truth||Prediction|+|Ground $\text{Dice}=\frac{2\times\left|\text{Prediction}\cap\text{Ground Truth}\right|}{\left|\text{Prediction}\right|+\left|\text{Ground Truth}\right|}$ (6)
@@ 3.2 Segmentation of galaxies based on U-Net
After advanced segmentation techniques, the next key step was to extract relevant data from the FITS files using the resulting masks. This process is crucial to use our preprocessed galaxy images for further understanding and analysis. The result of this extraction can be seen in Fig. 8. This approach allows us to efficiently search existing galaxy databases and use the compiled data for a comprehensive analysis, which is detailed in the next section.
@@
In Fig. 9 we illustrate the whole workflow of efficient processing of galaxy image data to ensure reliable preparation, modeling, and post-processing, developed in this work. This workflow is fundamental to understanding a systematic approach to the analysis of astronomical images and provides a clear path from raw data acquisition to the production of usable scientific data. To ensure complete transparency of our process and facilitate reproducibility, we provide comprehensive documentation, including all codes used in our analyses, available on our GitHub repository [^2] https://github.com/MatFyzGalaxy/EdgeOnML.
Figure 9: Diagram outlining the complete workflow used in the study, illustrating the process of data preparation, modelling and post-processing. Starting with the pre-processing phase, we transform the FITs data into JPEG format and annotate it, followed by splitting it into labeled datasets for machine learning training. The modeling phase involves training the YOLOv5 detection model and the SCSS-net segmentation model to identify and segment galaxies from astronomical images. The post-processing phase focuses on generating segmentation masks and extracting galaxy parameters for further analysis, demonstrating an integrated approach from raw data to usable scientific outputs.
@@ 4 Properties of the selected sample
Table 3: Selected parameters of the detected sample, taken from SDSS, unless stated otherwise. The full version of the table is available at our Github repository https://github.com/MatFyzGalaxy/EdgeOnML. Columns: (1): Unique SDSS identifier, (2): Galaxy name from NED, (3) and (4): J2000 coordinates, (5): Petrosian radius in the $r$ band, (6)-(10): Petrosian flux in the respective bands, (7): Exponential fit $b/a$ in the $r$ band, (8): Redshift, (9): Confidence with which the galaxy was detected, given by the algorithm.
1237667734496084000 UGC 05535 153.939 18.946 16.652 14.914 14.047 13.617 13.315 0.477 0.022 0.86
1237665228376899651 2MASS J11544292+3332148 178.680 33.537 17.963 15.543 14.762 14.315 14.002 0.220 0.032 0.86
1237661069790675120 NGC 3419A 162.833 14.023 15.783 14.694 14.263 14.046 13.825 0.197 0.010 0.85
1237661358608613531 UGC 05531 153.929 43.993 16.899 14.912 14.148 13.735 13.370 0.319 0.017 0.85
1237664671645368370 2MASS J13364223+3444493 204.175 34.746 16.480 14.706 13.903 13.466 13.192 0.386 0.024 0.84
1237662224593977484 2MASS J10211082+3539315 155.295 35.660 23.218 21.888 20.781 20.105 19.666 0.295 - 0.84
1237655126083043481 NGC 4252 184.629 5.559 15.665 14.516 14.126 13.871 13.777 0.312 0.003 0.84
1237662524157526119 2MASS J12412516+1343485 190.356 13.730 15.849 14.670 14.048 13.726 13.472 0.230 0.020 0.84
1237667142867615779 SDSS J092440.07+250642.6 141.167 25.110 16.339 14.574 13.698 13.227 12.885 0.414 0.027 0.84
1237667322179616901 2MASS J11425211+2632287 175.718 26.541 17.259 15.278 14.324 13.828 13.503 0.412 0.028 0.83
In this section, we show some basic properties of the sample detected by our algorithm and show that in general they are of the same type of galaxies that we trained the algorithm on.
After segmentation and extraction of the galaxies from their fits files, we query their coordinates in the SDSS catalogue, retrieving their redshifts, magnitudes and $b/a$ ratio using the SciServer package. We also retrieve the names of the galaxies from the NED database using the astroquery package (Ginsburg et al., 2019). In Table 3 we show the parameters of the first ten galaxies from our sample. The full table is available at our Github repository as well.
@@
from Blanton (2006), where $Q_{0}=2.8$, $Q_{1}=-1.8$ and $z_{0}=0.1$. This formula was derived for the SDSS sample, with redshifs $0.150.05<z<$ 0.15, which is satisfied for our sample.
@@
In Fig. 11 we show an example of randomly selected galaxies with the respective confidence given by the algorithm. By inspecting the galaxies visually, we can see that the algorithm is rather conservative, and galaxies that are clearly edge-on or highly inclined are detected with confidence $0.7-0.9$. Galaxies with confidence $<0.5$ are mostly small and/or faint, therefore, their nature could be debated. The default confidence threshold for YOLOv5 is set to 0.25. This value determines the minimum confidence score required for the model to consider a detected object as valid. Adjusting this threshold can help balance between detecting all possible objects (sensitivity) and minimising FP (precision). After optimising on our validation dataset, we used an optimal confidence threshold of $0.3$ to achieve a better balance between precision and recall, ensuring more reliable detections in our specific context. Therefore, we provide all galaxies detected with confidence $>0.3$, but for further astrophysical analysis we recommend using galaxies with confidence at least $>0.5$.
@@
Looking closer at our sample of galaxies, in Fig. 12 we can see that our galaxies are nearby, with most of the redshifts being about $0.100.02<z<$ 0.10, which is expected for our sample and is in good agreement with the redshifts of the Galaxy Zoo sample that we used for training. In Fig. 13 we show the colour-magnitude diagram of our sample. We divide the magnitude into red, green and blue parts, based on criteria from Papastergis et al. (2013). Most of our galaxies are red, which may seem unexpected, since we are looking for spiral galaxies, but this result is in agreement with Makarov et al. (2022) who made a catalogue of $16,551$ edge-on galaxies created using the public DR2 data of the Pan-STARRS survey. They show that edge-on galaxies are redder than general samples of galaxies containing all inclinations, due to internal extinction of edge-on galaxies, which is also why these galaxies are usually excluded when calculating Tully-Fisher relationship (see for example Karachentsev et al., 2002).
@@
The minor to major axes ratio $b/a$, which can be seen in Fig. 14 is low for most detected galaxies, which is expected from edge-on galaxies and is in agreement with Makarov et al. (2022). We see that there are some outliers that seem thicker, which is due to several factors. Some galaxies detected by the algorithm are faint and/or blurry, especially those detected with lower confidence, therefore, they might have higher values of $b/a$. Moreover, our original sample of galaxies from Galaxy Zoo, which might have some annotation bias, has similar distribution of $b/a$, so our algorithm is detecting galaxies similar to what it was trained on. Another point is that some galaxies that have a relatively high value of $b/a$ in the SDSS catalogue appear as thin and edge-on or highly inclined. We illustrate this in Fig. 10, where we show some examples of galaxies with $0.390.35<(b/a)_{r}<$ 0.39, although they appear as thin and edge-on, therefore, the algorithm detected them correctly.
Makarov et al. (2022) also found that on average, red galaxies are thicker than blue ones. In order to corroborate this, we plot a normalised histogram of inverse galaxy thickness in Fig. 15 for the red and blue populations separately. We see that our red population reaches a bit higher values of thickness, however the maximum of both red and blue sample lies around the same value of $a/b$ and the difference between the samples is rather small. Moreover, the sample of blue galaxies is quite smaller than the red sample, which may bias a direct comparison, therefore, we cannot make a definitive conclusion about the colour-thickness relationship.
Figure 12: Distribution of redshifts of the sample.
Figure 13: Galaxy colour-magnitude diagram for our sample, separated into red, green and blue ares, using only galaxies for which the redshift is known. The colours were corrected for Galactic extinction and $K$ correction as described in Section 4. The red line is defined as $g-i=-0.0571(M_{r}+24)+1.2$, while the blue line is offset by $0.15$ mag (Papastergis et al., 2013)
Figure 14: Minor to major axes ratio as a function of absolute magnitude, using only galaxies for which the redshift is known.
Figure 15: Normalised histogram of the inverse thickness of the galaxies for red and blue galaxies. The galaxies were separated as in Fig. 13.

### v2

@@ 3.4 Evaluation of YOLO-based Detection Results
The results of the performance of the YOLOv5 model on the test set are summarized in Table 1. The table shows the number of images and its instances of edge-on galaxies, along with the key metrics: Precision and Recall. These results indicate that the model performs well in detecting edge-on galaxies, with a balanced trade-off between precision and recall.
@@
Table 1: Table with metrics calculated by YOLOv5 on testing dataset. One image may contain more than 1 instance of edge-on galaxy. Precision and Recall are calculated according to True Positive (TP), False Positive (FP) and False Negative (FN) values (see Table 2).
@@
To provide a comprehensive understanding of the performance of the YOLOv5 model, we also provide particular detection cases, where we can see the sources of errors more precisely. Table 2 summarizes the counts of true positives (TP), false positives (FP), and false negatives (FN), which are essential to estimate the detection capabilities of the model.
Table 2: Summary of True Positives (TP), False Positives (FP), and False Negatives (FN) identified by YOLOv5 detection model.
@@
Due to manual annotations by students, there is some inevitable annotation bias, e.g., not every time students annotated all instances or sometimes a wrong object was annotated. One of the advantages of deep learning techniques is their ability to overcome annotation bias issues and provide correct predictions even for cases with annotation mistakes. This is possible if annotations are generally correct for most cases, and the architecture is able to learn a good generalized prediction function. Hence, the real efficiency of the algorithm is even better and might be additionally estimated by the analysis of wrong cases.
According to our knowledge from the Galaxy Zoo project, there should be at least one edge-on galaxy in the image, but that does not rule out the possibility of detecting other galaxies. Initially, our annotators only flagged galaxies that they were confident of, but our algorithm successfully detected several others. This is illustrated in Fig. 5, where we present an example image that initially had one galaxy annotation from Zooniverse. As can be seen in Fig. 6, after applying a trained YOLOv5 model, an additional galaxy was detected. This demonstrates the capability of the model to identify galaxies that may have been missed in the initial annotations due to manual annotation bias.
@@
To verify these findings, we manually examined detections labeled false positives (FP) and false negatives (FN). Figures 7 and 8 show cases of false positives and false negatives, providing visual context for these detections.
@@
Through visual inspection, we found that a number of misclassifications are due to inaccuracies in the annotation. In more detail, at least $\sim 20\%$ of the FP cases are considered edge-on galaxies, which were not annotated by mistake, and therefore these cases might be relocated from FP to TP after such analysis. For another $\sim 20\%$ of galaxies we were not able to decide, based on visual inspection, what their inclination is, but some of them might be edge-on as well. Hence, we estimate that the precision is probably at least $\sim 85\%$ in comparison to $\sim 80\%$ measured for the original annotations. Although these numbers are only approximate, as they were obtained by visual inspection, which is prone to bias, they suggest that performance indicators may be underestimated and could potentially be higher with more accurate initial data. This observation highlights one of the key strengths of deep learning models mentioned before: the ability to learn the correct prediction function, even in the presence of partially imperfect annotations.
3.5 Evaluation of U-Net Segmentation Results
@@
Table 3: Selected parameters of the detected sample, queried from the SDSS database, unless stated otherwise. The full version of the table is available at our Github repository https://github.com/MatFyzGalaxy/EdgeOnML. Columns: (1): Unique SDSS identifier, (2): Galaxy name from NED, (3) and (4): J2000 coordinates, (5)-(9): Petrosian flux in the respective bands, (10): Exponential fit $b/a$ in the $r$ band, (11): Redshift, (12): Confidence with which the galaxy was detected, given by the algorithm.
1237663237128519791 UGC 02699 50.5042 -1.0535 18.281 16.143 15.209 14.738 14.324 0.275 0.0266 0.861
1237665369582928005 UGC 06539 173.3493 32.6037 21.547 21.176 21.645 22.773 22.827 0.999 - 0.859
1237648705119060162 FGC 1534 195.1273 0.4801 17.952 16.648 15.897 15.418 15.067 0.190 0.0682 0.840
1237667730200264899 UGC 04742 135.5952 14.5252 17.095 15.317 14.437 13.969 13.611 0.444 0.0301 0.839
1237661358611824658 CGCG 241-036 163.6793 46.5540 16.461 14.683 13.828 13.387 13.070 0.485 0.0222 0.837
1237655126084878404 VCC 1597 188.7607 5.4249 17.688 16.175 15.533 15.219 15.284 0.204 - 0.836
1237667734496084000 UGC 05535 153.9388 18.9461 16.652 14.914 14.047 13.617 13.315 0.477 0.0216 0.835
1237667322722844715 UGC 07959 191.9249 26.9816 16.183 14.344 13.510 13.096 12.789 0.473 0.0239 0.834
1237663785276670105 FGC 0096 12.7149 0.8513 17.906 16.511 16.104 15.839 15.817 0.123 0.0153 0.833
1237657606426263745 CGCG 208-032 129.4910 39.8962 17.073 15.478 14.642 14.208 13.823 0.426 0.0242 0.832
@@
After advanced segmentation techniques, the next key step was to extract relevant data from the FITS files using the resulting masks. This process is crucial to process our galaxy images for further understanding and analysis. The result of this extraction can be seen in Fig. 11. This approach allows us to efficiently search existing galaxy databases and use the compiled data for a comprehensive analysis, which is detailed in the next section.
To ensure complete transparency of our process and facilitate reproducibility, we provide comprehensive documentation, including all codes used in our analyses, available on our GitHub repository [^4] https://github.com/MatFyzGalaxy/EdgeOnML.
@@ 4 Properties of the selected sample
In this section, we show some basic properties of the sample detected by our algorithm and show that in general they are of the same type of galaxies that we trained the algorithm on. The finished algorithm was run on the full SDSS DR17 sample (Abdurro’uf et al., 2022), where we detected $171,946$ objects. However, as we discuss below, many of these are detected with a low confidence as they are faint and blurry, so in further analysis we only consider galaxies detected with $Confidence>0.7$, which results in about $\sim 12,000$ galaxies.
After segmentation and extraction of the galaxies from their fits files, we queried their coordinates in the SDSS catalog, retrieving their redshifts, magnitudes and $b/a$ ratio using the SciServer package. We also retrieved the names of the galaxies from the NED database using the astroquery package (Ginsburg et al., 2019a). In Table 3 we show the parameters of the first ten galaxies from our sample. The full table is available at our Github repository as well.
@@
from Blanton (2006), where $Q_{0}=2.8$, $Q_{1}=-1.8$ and $z_{0}=0.1$. This formula was derived for the SDSS sample, with redshifts $0.150.05<z<$ 0.15, which is satisfied for our sample.
@@
In Fig. 12 we show an example of randomly selected galaxies with the respective confidence given by the algorithm. By inspecting the galaxies visually, we can see that the algorithm is rather conservative, and galaxies that are clearly edge-on or highly inclined are detected with $0.90.7<Confidence<$ 0.9. Galaxies with $Confidence<0.5$ are mostly small and/or faint, therefore, their nature could be debated. The default confidence threshold for YOLOv5 is set to $0.25$. This value determines the minimum confidence score required for the model to consider a detected object as valid. Adjusting this threshold can help balance between detecting all possible objects (sensitivity) and minimizing false positives (precision). After optimising on our validation dataset, we used an optimal confidence threshold of $0.3$ to achieve a better balance between precision and recall, ensuring more reliable detections in our specific context. Therefore, we provide all galaxies detected with $Confidence>0.3$, but for further astrophysical analysis, we recommend using galaxies with at least $Confidence>0.5$ or higher. In the following analysis, we only consider galaxies with $Confidence>0.7$.
Figure 13: Distribution of redshifts of the detected sample with $Confidence>0.7$ are considered.
@@
Looking closer at our sample of galaxies, in Fig. 13 we can see that our galaxies are nearby, with most of the redshifts being about $0.100.02<z<$ 0.10, which is expected for our sample and is in good agreement with the redshifts of the Galaxy Zoo sample that we used for training. In Fig. 14 we show the colour-magnitude diagram (CMD) of our sample compared to the CMD of the training sample obtained from Galaxy Zoo 2 as described in Section 2. We divide the magnitude into red, green and blue parts, based on criteria from Papastergis et al. (2013). Most of our galaxies are red, which may seem unexpected, since we are looking for spiral galaxies, but this result is in agreement with Makarov et al. (2022) who made a catalog of $16,551$ edge-on galaxies created using the public DR2 data from the Pan-STARRS survey. They show that edge-on galaxies are redder than general samples of galaxies containing all inclinations, due to internal extinction of edge-on galaxies, which is also why these galaxies are usually excluded when calculating Tully-Fisher relationship (see for example Karachentsev et al., 2002). Moreover, there is excellent agreement between our detected sample and the training sample, showing that our algorithm reproduces the properties of the original sample.
The minor to major axes ratio $b/a$, which can be seen in Fig. 15 is low for most detected galaxies, which is expected from edge-on galaxies and is in agreement with Makarov et al. (2022). We see that there are some outliers that seem thicker; however, when we compare our sample with the training sample from Galaxy Zoo, we see that this sample has a similar trend of $b/a$. Therefore, our algorithm is correctly detecting galaxies similar to what it was trained on. Moreover, on average the detected sample has lower values of $b/a$ than the training sample, so our algorithm shows some improvement compared to the training sample, which may suffer from an annotation bias. Another point is that some galaxies that have a relatively high value of $b/a$ in the SDSS catalog appear as thin and edge-on or highly inclined. We illustrate this in Fig. 16, where we show some examples of galaxies with $0.390.35<(b/a)_{r}<$ 0.39, although they appear as thin and edge-on, therefore, the algorithm detected them correctly.
Makarov et al. (2022) also found that on average, red galaxies are thicker than blue ones. In order to corroborate this, we plot a normalized histogram of inverse galaxy thickness in Fig. 17 for the red and blue populations separately. We confirm this finding, showing that red galaxies tend to be more thick, while galaxies with lower thickness are more likely to be blue.
We also compared our detected sample to the catalog of Bizyaev et al. (2014) who compiled a catalog of $5,747$ edge-on galaxies from the SDSS DR7 supplemented by data from other surveys (Revised Flat Galaxy Catalog (Karachentsev et al., 1999), RC3 (de Vaucouleurs et al., 1991), EFIGI (Baillard et al., 2011), and Galaxy Zoo (Lintott et al., 2011)), selected by visual classification. A simple cross-match of our samples reveals that we detected approximately $3,500$ identical galaxies. Considering that we use two different SDSS data releases and that Bizyaev et al. (2014) also added galaxies from a few different catalogs, we find this agreement reasonable. Moreover, as we showed, we found a good agreement of the properties of our sample with previous studies.
Figure 14: Galaxy colour-magnitude diagram separated into red, green, and blue areas, using only galaxies for which the redshift is known. The colours were corrected for Galactic extinction and $K$ correction as described in Section 4. The red line is defined as $g-i=-0.0571(M_{r}+24)+1.2$, while the blue line is offset by 0.15 mag (Papastergis et al., 2013). Left: Sample of detected galaxies with $Confidence>0.7$. Right: Galaxy Zoo Sample used for training.
Figure 15: Minor to major axes ratio as a function of absolute magnitude, using only galaxies for which the redshift is known. Left: Sample of detected galaxies with $Confidence>0.7$. Right: Galaxy Zoo Sample used for training.
@@
Figure 17: Normalized histogram of the inverse thickness of the detected galaxies for red and blue galaxies. Only galaxies with $Confidence>0.7$ are considered. The galaxies were separated as in Fig. 14.